OpenStudy (tate_langdon):

Can someone check my work and show work Simplify the following expression. 1/8 + 1/7 A. 1/28 B. 2/15 C. 15/56 D. 15/122

OpenStudy (tate_langdon):

I say the answer is C

OpenStudy (tate_langdon):

@sunnnystrong

OpenStudy (sunnnystrong):

\[\frac{ 1 }{ 8 }+\frac{ 1 }{ 7 }=\frac{ 7 }{ 56 }+\frac{ 8 }{ 56 }=\frac{ 15 }{ 56 }\] Yep looks good :D

OpenStudy (tate_langdon):

Thanks

OpenStudy (sunnnystrong):

NP :D

OpenStudy (tate_langdon):

Yeah Tack sa mycket!!!